What is the y-intercept for y+2=-2?
Answer by xcentaur(357)

y+2=-2
i can't see any 'x' here,so if this is your correct equation,then this is how you'd go about it:
y+2=-2
y=-2-2
y=-4


This means that the y-coordinate for the line would be '-4' for all values of x.
Your points would become:
(0,-4)(1,-4)(2,-4) and so on


As you can see,the intercept of this line would be (0,-4)
Hope this helps,
Prabhat